OpenStudy (anonymous):

Solve the following equation: 4-e^x = 3e^(x+2) with the answer ln(4/(3e^(2) - 1).

OpenStudy (jamesj):

e^(x+2) = e^2.e^x Now use that to find an expression e^x = something. Then take the log of both sides.

OpenStudy (jamesj):

Making sense?

OpenStudy (jamesj):

(and btw, I think there's a sign wrong in your answer.)

OpenStudy (anonymous):

I get the concept, but I can't get it to work...

OpenStudy (jamesj):

4-e^x = 3e^(x+2) 4 - e^x = 3.e^2.e^x 4 = 3e^2.e^x + e^x (3e^2 + 1)e^x = 4 e^x = 4/(3e^2 +1)

OpenStudy (anonymous):

Ah, fantastic - I got to the second step but couldn't see what to do next! Obvious now, of course... Thank you!